# Dynamic thermalization on noisy quantum hardware

### v1

@@ Appendix A Variational circuit recompilation
Appendix B Error mitigation
@@
In addition to the qubits on which the CNOT gate is applied, a noisy CNOT gate may also lead to coherent errors on neighboring qubits. In our case, this can be formalized by extending the noise channel associated with a given CNOT gate to all qubits in the register. This noise, known as spillover crosstalk noise [60], can also be made purely incoherent by extending the RC method to neighboring qubits. In addition, since a priori no logical operation is being applied on these other qubits, the RC procedure can be improved by applying extra random $\pi/2$ rotations on top of the Pauli matrices, as prescribed in [36]. One can show that by doing this, the effective incoherent noise channel obtained after averaging over crosstalk RC circuits is a combination of a local Pauli noise channel for the active qubits of the CNOT gates, a local depolarizing noise channel acting on the neighboring qubits, and a global depolarizing noise channel acting uniformly on all qubits. In our quantum simulation, we run $100$ crosstalk RC circuits for every realization of the quench protocol, and average the read-out observables, effectively realizing the noise channel described above. We have checked for this number of circuits that the crosstalk RC averages has clearly converged to its infinite sampling value. Using our knowledge of the structure of the effective incoherent noise channel acting on our qubits, we now implement an error mitigation strategy. Below, we describe how the effect of the global depolarizing noise channel can be removed by a simple fit. Then, we leverage the errors due to the local Pauli and depolarizing noise channels using zero-noise extrapolation (ZNE), as described in the main text.
@@
where $\rho$ the density matrix, $f$ is the fidelity of a CNOT gate and $\mathbb{1}$ the identity matrix of size $2^{N}\!$. The rightmost term, which models the output of a global depolarizing error occurring, is the maximally mixed state with $N$ qubits. If such an error occurs on any CNOT gate during the execution of the circuit, then the qubits will remain in this state, as it is not affected by the action of subsequent unitary operations. Thus, it is straightforward to generalize Eq. (12) to the case of a circuit with $N_{\text{CNOT}}$ CNOT gates by replacing the fidelity of a single CNOT gate $f$ with a circuit fidelity $f^{N_{\rm CNOT}}$. Further averaging the quantum simulation data for occupation probabilities over $100$ realizations of the quench protocol, we interpret the noisy quantum computer’s output within the model (12), leading to the expression (11) for noisy average occupations in the main text.

### v2

@@ Appendix B Variational circuit recompilation
Appendix C Error mitigation
@@
In addition to the qubits on which the CNOT gate is applied, a noisy CNOT gate may also lead to coherent errors on neighboring qubits. In our case, this can be formalized by extending the noise channel associated with a given CNOT gate to all qubits in the register. This noise, known as spillover crosstalk noise [75], can also be made purely incoherent by extending the RC method to neighboring qubits. In addition, since a priori no logical operation is being applied on these other qubits, the RC procedure can be improved by applying extra random $\pi/2$ rotations on top of the Pauli matrices, as prescribed in [42]. One can show that by doing this, the effective incoherent noise channel obtained after averaging over crosstalk RC circuits is a combination of a local Pauli noise channel for the active qubits of the CNOT gates, a local depolarizing noise channel acting on the neighboring qubits, and a global depolarizing noise channel acting uniformly on all qubits. In our quantum simulation, we run $100$ crosstalk RC circuits for every realization of the quench protocol, and average the read-out observables, effectively realizing the noise channel described above. We have checked for this number of circuits that the crosstalk RC averages has clearly converged to its infinite sampling value. Using our knowledge of the structure of the effective incoherent noise channel acting on our qubits, we now implement an error mitigation strategy.
@@
where $\rho$ the density matrix, $f$ is the fidelity of a CNOT gate and $\mathbb{1}$ the identity matrix of size $2^{N}\!$. The rightmost term, which models the output of a global depolarizing error occurring, is the maximally mixed state with $N$ qubits. If such an error occurs on any CNOT gate during the execution of the circuit, then the qubits will remain in this state, as it is not affected by the action of subsequent unitary operations. Thus, it is straightforward to generalize Eq. (12) to the case of a circuit with $N_{\text{CNOT}}$ CNOT gates by replacing the fidelity of a single CNOT gate $f$ with a circuit fidelity $f^{N_{\rm CNOT}}$. Further averaging the quantum simulation data for occupation probabilities over $100$ realizations of the quench protocol, we interpret the noisy quantum computer’s output within the model (12), leading to the expression (11) for noisy average occupations in the main text. Then, we leverage the errors due to the local Pauli and depolarizing noise channels using zero-noise extrapolation (ZNE).